# Chain-of-Thought Prompting Obscures Hallucination Cues in Large Language Models: An Empirical Evaluation

### v1

@@ 5.2 Comparison of Score Distributions for Hallucination Detection
We use the Kolmogorov–Smirnov (K–S) test to evaluate whether these distributions differ significantly across conditions. As shown in Figure 4, both scores undergo statistically significant shifts (p < 0.01), indicating that CoT prompts induce meaningful changes in the LLM’s internal representations.
@@
For instruction-tuned LLMs like LLaMA, the scores are relatively dispersed without CoT, suggesting that the hallucination detection methods can effectively differentiate between hallucinated and faithful responses. However, with CoT prompting, the scores become tightly clustered, implying that the detector may collapse to predicting a single class—thereby losing discriminative capability.
In contrast, reasoning-oriented LLMs such as DeepSeek exhibit concentrated score distributions even without CoT, suggesting that internal signals are less indicative of hallucination. CoT prompting still significantly alters the distribution, suggesting that explicit reasoning substantially reshapes the LLM’s internal state—even when the LLM is already strong in reasoning tasks.
Table 3: Number of experiments exhibiting AUROC decline across different LLMs and datasets.
@@ 5.3 Performance Comparison of Hallucination Detection Methods
We conduct a comprehensive evaluation of CoT prompting affects the performance of hallucination detection methods. Our experiments span three CoT prompting methods, four benchmark datasets, four LLMs, two hallucination annotation protocols (i.e., methods for establishing ground truth), and eight representative detection approaches, yielding a total of 768 experimental configurations. Table 3 summarizes the proportion of settings in which the AUROC score declined after applying CoT prompting. Overall, more than half of the setups (440 out of 768) exhibit performance degradation, suggesting that CoT prompting exerts a broad negative impact on hallucination detection. Notably, we treat TruthfulQA separately due to its distinct evaluation protocol.
@@
Our findings reveal that CoT prompting generally impairs the effectiveness of hallucination detection, with the degree of impact varying by method. Self-evaluation-based approaches (e.g., SelfCheckGPT-Prompt and Verbalized Certainty) are most adversely affected. These methods rely heavily on confidence signals present in the LLM’s output, which are easily distorted by CoT-style reasoning. CoT prompting often causes LLMs to generate additional justificatory content that obscures factual inaccuracies. For instance, in response to the TriviaQA question What was Walter Matthau’s first movie?, an LLM under LtM prompting may include phrases like Based on reliable sources such as IMDb and Wikipedia…. Although the answer is incorrect, the Verbalized Certainty score may increase from 0.2 to 1.0, illustrating how CoT prompting can mask hallucinations.
@@
By contrast, consistency-based methods (e.g., EigenScore and SelfCheckGPT-NLI) show greater robustness to CoT prompting. These methods leverage multi-sample reasoning to identify inconsistencies, which remain informative even in the presence of CoT prompts. In the above example, consistent sampling occasionally yields the correct answer The Kentuckian (EigenScore: 3/15; SelfCheckGPT-NLI: 1/20), albeit infrequently. While often criticized for high computational overhead, these approaches demonstrate strong resilience and reliability in reasoning-heavy scenarios.
@@
A further notable observation is that although advanced hallucination detection methods typically outperform simple perplexity-based baselines under standard conditions, this advantage diminishes with CoT prompting. Without CoT, perplexity-based methods consistently underperform, aligning with prior studies (Chen et al., 2024a; Sriramanan et al., 2024). However, when CoT prompts are introduced, many detection methods fail to surpass the perplexity baseline, and some perform worse. This finding underscores a critical limitation of current detection approaches: they are not yet equipped to handle the structural and discourse-level shifts induced by explicit reasoning, ultimately reducing their ability to detect hallucinations effectively.
@@
TruthfulQA. As shown in Table 4, CoT prompting exhibits inconsistent effects on enhancing the truthfulness and informativeness of responses in the TruthfulQA task across different scenarios (see Appendix G for details). Meanwhile, the negative impact of CoT prompting on hallucination detection methods varies significantly across different LLMs. Specifically, Llama shows a more pronounced performance drop in informativeness, Mistral experiences a greater decline in truthfulness, and DeepSeek shows a decline in both dimensions. This variation likely reflects differing response generation tendencies among LLMs. For example, after applying CoT prompting, Llama tends to emphasize the reasoning process and often includes redundant phrases such as Based on the above steps… in its answers. While such content highlights the reasoning chain, it may obscure factual inaccuracies or introduce misleading elements, thereby interfering with the judgment of hallucination detection methods.
@@
Table 4: Number of experiments in which the AUROC value decreases. For each experiment, AUROC is calculated by comparing the predicted hallucination detection scores with either informativeness (I.) or truthfulness (T.); each cell shows the number of experiments (three using CoT prompting and one without) where AUROC decreases relative to a baseline. The first two columns are results for Llama-3.1-8B-Instruct, the middle two for Mistral-7B-Instruct-v0.3, and the last two for DeepSeek-R1-Distill-Llama-8B.
@@ 5.4 Confidence in Hallucination Detection Methods
We further employ Expected Calibration Error (ECE) to evaluate the alignment between the hallucination probabilities predicted by the detection methods and the actual correctness of the hallucination detection, thereby assessing the calibration of model confidence. The results indicate that CoT prompting leads to an increase in ECE across most experimental settings, suggesting that the confidence of detection methods becomes less reliable. Detailed results are provided in Appendix G.
@@ 6 Discussion and Conclusion
In this study, we systematically investigate the impact of CoT prompting on hallucination detection in LLMs. We evaluate several representative detection methods and find that while CoT prompting improves text-level and semantic accuracy across most datasets, it simultaneously undermines detection performance. More specifically, CoT prompting leads to more concentrated detection score distributions, reduced classification accuracy, weaker alignment with ground truth, and poorer confidence calibration. The sensitivity of impact varies by method: self-evaluation-based approaches are most affected, whereas consistency-based methods exhibit greater robustness. We also observe that CoT prompting also diminishes, and occasionally eliminates, the performance advantage of advanced detectors over a simple perplexity baseline.
These findings suggest that CoT prompting can obscure hallucination cues and compromise existing detection systems, highlighting the need for more robust approaches tailored to reasoning-oriented LLM.
@@ B.1 Large Language Models and Hallucinations
Large language models demonstrate remarkable generalization capabilities, leading to their widespread adoption across a variety of tasks (Zhao et al., 2024; Qin et al., 2024; Yan et al., 2024). They exhibit strong performance, especially in question-answering tasks (Zhao et al., 2024; Yang et al., 2023; Hadi et al., 2024). However, despite their effectiveness, LLMs are not infallible, and a significant challenge that persists is the phenomenon of hallucinations (Ji et al., 2023; Yang et al., 2023; Hadi et al., 2024).
Hallucinations in LLMs typically manifest when the model generates fabricated or inaccurate information in an attempt to answer a question (Yang et al., 2024b; Tonmoy et al., 2024). These hallucinations can be classified into three categories: Input-Conflicting, Context-Conflicting, and Fact-Conflicting hallucinations (Zhang et al., 2023). Alternatively, they are often categorized into two broad types: factuality hallucination and faithfulness hallucination (Huang et al., 2024). Factuality hallucinations arise when the LLM’s output conflicts with established world knowledge. Specifically, these hallucinations arise when the input prompt lacks relevant factual information. In such cases, the LLM relies on its internal parameters and memory (Mallen et al., 2023), resulting in outputs that contradict factual reality.
@@ B.2 CoT prompting
Chain-of-Thought (CoT) prompting is a technique designed to enhance the performance of LLMs by guiding them to generate intermediate reasoning steps prior to producing a final answer (Wei et al., 2022). Rather than generating a direct one-step response, CoT prompts the LLM to engage in a structured reasoning process, thereby reducing errors and hallucinations that arise from premature conclusions. By leveraging this step-by-step reasoning approach, CoT demonstrates significant improvements in both accuracy and reliability.
To reduce the manual effort required to provide examples and to better leverage LLMs’ prior knowledge, Self-generated Few-shot CoT is introduced (Yasunaga et al., 2024). This approach allows LLMs to autonomously construct relevant examples and knowledge, enabling adaptive reasoning. Remarkably, studies demonstrate that performance improvements associated with reasoning can occur even without explicitly providing examples (Kojima et al., 2024). By simply appending a phrase such as “let’s think step by step” to the prompt — a technique known as Zero-shot CoT — LLMs are prompted to generate reasoning steps. This appended phrase (referred to as a “magic phrase” (Stechly et al., 2024)) is identified as a key trigger for enabling reasoning without explicit demonstrations.
@@ B.3 Hallucination Detection
Consistency-based methods, such as SelfCheckGPT, operate under the assumption that if an LLM has accurate knowledge and confidence in its answers, it will generate semantically consistent outputs across multiple sampling attempts (Wang et al., 2023; Manakul et al., 2023). This approach is particularly relevant in factual QA tasks, where semantic consistency among sampled answers serves as an indicator of factuality. Inconsistent outputs, on the other hand, suggest the presence of hallucinations.
@@ Appendix E Details on the Evaluation Framework
Distribution Consistency Test: We use statistical methods, such as the Kolmogorov-Smirnov test, to compare the distribution of hallucination scores with and without CoT prompting. This analysis determines whether the introduction of CoT prompting significantly alters the distribution of hallucination scores. It also provides an intuitive measure of how CoT prompting influences hallucination detection.
Classification Performance: We treat correctness as the ground truth (binary: correct or incorrect) and hallucination scores as predictions. We then use AUROC to quantify the performance of hallucination detection methods before and after applying CoT prompting. This metric quantifies the degree to which CoT prompting influences the effectiveness of hallucination detection methods.
Monotonic Relationship: Using Spearman’s rank correlation coefficient, we evaluate the monotonic relationship between hallucination scores and correctness values. Here, correctness is treated as a continuous variable rather than a binary label. This metric shows whether hallucination detection scores are consistent with correctness under CoT prompting.
Alignment: We use Expected Calibration Error (ECE) to assess how well hallucination scores align with correctness. By treating correctness as the ground truth and hallucination scores as predictions, this metric evaluates whether hallucination detection methods effectively reflect the degree of correctness. This approach goes beyond simply distinguishing between correct and incorrect responses. This is particularly important as some answers may not be strictly correct or incorrect, and well-calibrated hallucination scores should reflect the nuances of correctness.
@@ F.2 Main Experiment
[…]
We validate our automated metrics by manually annotating 150 samples with 2 annotators, achieving a Cohen’s Kappa of 0.863. Spearman’s $\rho$ between our automated scores and human annotations is 0.690 for Exact Match, 0.656 for ROUGE-L F1, and 0.683 for semantic similarity, confirming that our automated evaluations reliably reflect human judgment.
@@
TruthfulQA offers a training set with 6.9k examples, which we use to fine-tune the GPT-3-6.7B model for evaluation purposes. Since OpenAI’s API no longer provides access to this model, we instead fine-tune the Llama-3.1-8B [^4] https://huggingface.co/meta-llama/Llama-3.1-8B model using LoRA (Hu et al., 2022) to compute the Truthfulness and Informativeness scores. We conduct fine-tuning on the LLaMA-Factory [^5] https://github.com/hiyouga/LLaMA-Factory platform with the following hyperparameters: learning rate = 5.0e-05, batch size = 32, epochs = 5, lora_rank = 8 and lora_alpha = 16.
The evaluation metrics are defined in Section 4.1. For Expected Calibration Error (ECE), we apply a normalization process to ensure predictions fall within the $[0,1]$ range. Specifically, we scale the data within the $3\%-97\%$ range to $[0,1]$, while outliers beyond this range are clipped to 00 or $1$. This normalization avoids distortions caused by extreme values.

### v3

@@ 5.2 Comparison of Score Distributions for Hallucination Detection
We use the Kolmogorov-Smirnov (K-S) test to evaluate whether these score distributions differ significantly between conditions with and without CoT prompting. As shown in Figure 4, both scores exhibit statistically significant shifts ($p<0.01$), indicating that CoT prompting induces substantial changes in the LLM’s internal representations.
@@
For instruction-tuned LLMs such as LLaMA, the scores are relatively dispersed in the absence of CoT prompting, suggesting that hallucination detection methods can effectively differentiate between hallucinated and faithful responses. However, with CoT prompting, the scores become tightly clustered. This implies that the detector may collapse to predicting a single class, thereby losing its discriminative capability.
In contrast, reasoning-oriented LLMs such as DeepSeek exhibit concentrated score distributions even without CoT, suggesting that internal signals are less indicative of hallucination. CoT prompting still significantly alters the distribution, suggesting that explicit reasoning substantially reshapes the LLM’s internal state, even when the LLM is already strong in reasoning tasks.
Table 3: Number of experimental configurations in which the AUROC for hallucination detection decreases after applying CoT prompting, relative to the base scenario. Each cell can contain up to 12 configurations, corresponding to 3 CoT variants $\times$ 4 LLMs.
@@ 5.3 Performance Comparison of Hallucination Detection Methods
We conduct a comprehensive evaluation of how CoT prompting affects the performance of hallucination detection methods. Our experiments span four benchmark datasets, four LLMs, two hallucination annotation protocols (i.e., Textual or Semantic methods for establishing ground truth), and eight representative detection approaches. For each combination, we evaluate performance under four prompting scenarios: one base scenario (without CoT) and three distinct CoT prompting methods (CoT, LtM, and MRPP). This results in a total of 768 experimental configurations (4 benchmark datasets $\times$ 4 LLMs $\times$ 2 hallucination protocols $\times$ 8 detection methods $\times$ 3 CoT prompting methods). For each configuration, the baseline performance without CoT prompting serves as the point of comparison.
Table 3 summarizes, grouped by LLM and dataset, the number of experimental configurations in which a detection method’s AUROC score declined under a CoT prompting scenario compared to the base scenario. Each cell in the table has a maximum value of 12, representing the 4 models evaluated under the 3 CoT methods. Overall, more than half of the configurations (465 out of 768) exhibit performance degradation, suggesting that CoT prompting exerts a broad and substantial negative impact on hallucination detection.
Notably, TruthfulQA is treated separately due to its distinct evaluation protocol. To further assess the generality of our findings, we also evaluate a model-training-based method, MIND (Su et al., 2024). We observe that it likewise suffers degraded hallucination detection performance under CoT prompting (see Appendix G for details).
@@
Our findings reveal that CoT prompting generally impairs the effectiveness of hallucination detection, with the degree of impact varying by method. Self-evaluation-based approaches (e.g., SelfCheckGPT-Prompt and Verbalized Certainty) are most adversely affected. These methods rely heavily on confidence signals present in the LLM’s output. However, these signals are easily distorted by CoT-style reasoning. As a result of CoT prompting, LLMs often generate additional justificatory content, which can obscure factual inaccuracies. For instance, in response to the TriviaQA question What was Walter Matthau’s first movie?, an LLM under LtM prompting may include phrases like Based on reliable sources such as IMDb and Wikipedia…. Although the answer is incorrect, the Verbalized Certainty score may increase from 0.2 to 1.0, illustrating how CoT prompting can mask hallucinations.
@@
By contrast, consistency-based methods (e.g., EigenScore and SelfCheckGPT-NLI) show greater robustness to CoT prompting. These methods leverage multi-sample reasoning to identify inconsistencies, which remain informative even in the presence of CoT prompts. In the above example, consistent sampling occasionally yields the correct answer The Kentuckian (EigenScore: 3/15; SelfCheckGPT-NLI: 1/20), albeit infrequently. Although these approaches are often criticized for their high computational overhead, they demonstrate strong resilience and reliability in scenarios that require extensive reasoning.
@@
Another notable observation is that, although advanced hallucination detection methods typically outperform simple perplexity-based baselines under standard conditions, this advantage diminishes when CoT prompting is applied. Without CoT, perplexity-based methods consistently underperform, aligning with prior studies (Chen et al., 2024a; Sriramanan et al., 2024; Chen et al., 2025). However, when CoT prompts are introduced, many detection methods fail to surpass the perplexity baseline, and some perform worse. This finding underscores a critical limitation of current detection approaches: they are not designed to handle the structural and discourse-level shifts induced by explicit reasoning. As a result, their effectiveness at detecting hallucinations is diminished.
@@
TruthfulQA. As shown in Table 4, CoT prompting has inconsistent effects on enhancing both the truthfulness and informativeness of responses in the TruthfulQA task across different scenarios. Among the 144 experimental settings, 81 exhibit performance degradation (see Appendix G for details). Meanwhile, the negative impact of CoT prompting on hallucination detection methods varies significantly among different LLMs. Specifically, Llama exhibits a more pronounced drop in informativeness, Mistral experiences a greater decline in truthfulness, and DeepSeek shows a decline in both dimensions. This variation likely reflects differences in response generation tendencies among the various LLMs. For example, following CoT prompting, Llama tends to emphasize the reasoning process and frequently includes redundant phrases, such as Based on the above steps…, in its answers. Although such content highlights the reasoning chain, it may obscure factual inaccuracies or introduce misleading elements. This can interfere with the effectiveness of hallucination detection methods. To mitigate potential biases introduced by the LLaMA-based judge model, we additionally fine-tuned the evaluation model using Qwen. The evaluation results based on the Qwen judge model (see Appendix G for details) are consistent with our previous findings, confirming the robustness of our conclusions.
@@
Table 4: Number of experiments of TruthfulQA in which the AUROC value decreases. For each experiment, AUROC is calculated by comparing the predicted hallucination detection scores with either informativeness (I.) or truthfulness (T.); each cell shows the number of experiments (three using CoT prompting and one without) where AUROC decreases relative to a baseline. The first two columns are results for Llama-3.1-8B-Instruct, the middle two for Mistral-7B-Instruct-v0.3, and the last two for DeepSeek-R1-Distill-Llama-8B.
@@ 5.4 Confidence in Hallucination Detection Methods
We further employ Expected Calibration Error (ECE) to evaluate the alignment between the hallucination probabilities predicted by the detection methods and the actual correctness of the hallucination detection. This allows us to assess the calibration of model confidence. The results indicate that CoT prompting leads to an increase in ECE across most experimental settings, suggesting that the confidence of detection methods becomes less reliable. Detailed results are provided in Appendix G.
@@ 6 Discussion and Conclusion
In this study, we systematically investigate the impact of CoT prompting on hallucination detection in LLMs. We evaluate several representative detection methods and find that, although CoT prompting improves text-level and semantic accuracy across most datasets, it simultaneously undermines hallucination detection performance. More specifically, CoT prompting results in more concentrated detection score distributions, lower classification accuracy, weaker alignment with ground truth, and poorer confidence calibration. The sensitivity of this impact varies according to the detection method: self-evaluation-based approaches are the most affected, whereas consistency-based methods show greater robustness. We further observe that CoT prompting diminishes and occasionally eliminates the performance advantage of advanced detectors over a simple perplexity baseline.
These findings suggest that CoT prompting can obscure hallucination cues and compromise the effectiveness of existing detection systems, highlighting the need to develop more robust approaches specifically tailored to reasoning-oriented LLMs.
@@
Acknowledgment
This work is supported by the National Natural Science Foundation of China (72204087), the Shanghai Planning Office of Philosophy and Social Science Youth Project (2022ETQ001), the Chenguang Program of Shanghai Education Development Foundation and Shanghai Municipal Education Commission (23CGA28), the Shanghai Pujiang Program (23PJC030), Young Elite Scientists Sponsorship Program by CAST (YESS20240562), and the Fundamental Research Funds for the Central Universities, China. We also appreciate the constructive comments from the anonymous reviewers.
@@ B.1 Large Language Models and Hallucinations
Large language models demonstrate remarkable generalization capabilities, leading to their widespread adoption across a variety of tasks (Zhao et al., 2024; Qin et al., 2024; Yan et al., 2024; Huang and He, 2025). They exhibit strong performance, especially in question-answering tasks (Zhao et al., 2024; Yang et al., 2023; Hadi et al., 2024). However, despite their effectiveness, LLMs are not infallible, and a significant challenge that persists is the phenomenon of hallucinations (Ji et al., 2023; Hadi et al., 2024; Wang et al., 2025).
Hallucinations in LLMs typically manifest when the model generates fabricated or inaccurate information in an attempt to answer a question (Yang et al., 2024b; Tonmoy et al., 2024). These hallucinations can be classified into three categories: Input-Conflicting, Context-Conflicting, and Fact-Conflicting hallucinations (Zhang et al., 2023). Alternatively, they are often categorized into two broad types: factuality hallucination and faithfulness hallucination (Huang et al., 2024). Factuality hallucinations arise when the LLM’s output conflicts with established world knowledge. Specifically, these hallucinations arise when the input prompt lacks relevant factual information (He et al., 2025). In such cases, the LLM relies on its internal parameters and memory (Mallen et al., 2023), resulting in outputs that contradict factual reality.
@@ B.2 CoT prompting
Chain-of-Thought (CoT) prompting is a technique designed to enhance the performance of LLMs by guiding them to generate intermediate reasoning steps prior to producing a final answer (Wei et al., 2022). Rather than generating a direct one-step response, CoT prompts the LLM to engage in a structured reasoning process, thereby reducing errors and hallucinations that arise from premature conclusions. By leveraging this step-by-step reasoning approach, CoT demonstrates significant improvements in both accuracy and reliability (Yao et al., 2025).
To reduce the manual effort required to provide examples and to better leverage LLMs’ prior knowledge, Self-generated Few-shot CoT is introduced (Yasunaga et al., 2024). This approach allows LLMs to autonomously construct relevant examples and knowledge, enabling adaptive reasoning. Remarkably, studies demonstrate that performance improvements associated with reasoning can occur even without explicitly providing examples (Kojima et al., 2024). By simply appending a phrase such as “let’s think step by step” to the prompt, a technique known as Zero-shot CoT, LLMs are prompted to generate reasoning steps. This appended phrase (referred to as a “magic phrase” (Stechly et al., 2024)) is identified as a key trigger for enabling reasoning without explicit demonstrations.
@@ B.3 Hallucination Detection
Consistency-based methods, such as SelfCheckGPT, operate under the assumption that if an LLM has accurate knowledge and confidence in its answers, it will generate semantically consistent outputs across multiple sampling attempts (Wang et al., 2023; Manakul et al., 2023; Ren et al., 2025). This approach is particularly relevant in factual QA tasks, where semantic consistency among sampled answers serves as an indicator of factuality. Inconsistent outputs, on the other hand, suggest the presence of hallucinations.
@@ Appendix E Details on the Evaluation Framework
Distribution Consistency Test. We use statistical methods, such as the Kolmogorov-Smirnov test, to compare the distribution of hallucination scores with and without CoT prompting. This analysis determines whether the introduction of CoT prompting significantly alters the distribution of hallucination scores. It also provides an intuitive measure of how CoT prompting influences hallucination detection.
Classification Performance. We treat correctness as the ground truth (binary: correct or incorrect) and hallucination scores as predictions. We then use AUROC to quantify the performance of hallucination detection methods before and after applying CoT prompting. This metric quantifies the degree to which CoT prompting influences the effectiveness of hallucination detection methods.
Monotonic Relationship. Using Spearman’s rank correlation coefficient, we evaluate the monotonic relationship between hallucination scores and correctness values. Here, correctness is treated as a continuous variable rather than a binary label. This metric shows whether hallucination detection scores are consistent with correctness under CoT prompting.
Alignment. We use Expected Calibration Error (ECE) to assess how well hallucination scores align with correctness. By treating correctness as the ground truth and hallucination scores as predictions, this metric evaluates whether hallucination detection methods effectively reflect the degree of correctness. This approach goes beyond simply distinguishing between correct and incorrect responses. This is particularly important as some answers may not be strictly correct or incorrect, and well-calibrated hallucination scores should reflect the nuances of correctness.
@@ F.2 Main Experiment
We conduct all experiments using four open-source instruction-tuned LLMs: Llama-3.1-8B-Instruct [^2] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct: A lightweight yet capable language model fine-tuned for instruction-following tasks, balancing efficiency and performance in resource-constrained environments (Dubey et al., 2024). Mistral-7B-v0.3-Instruct [^3] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3: A compact model optimized for competitive reasoning and low-latency inference, trained via advanced instruction tuning (Jiang et al., 2023). Llama-3.1-70B-Instruct [^4] https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct: A 70-billion-parameter Transformer-based model designed for high-precision instruction alignment, excelling in complex reasoning and human-preference tasks across diverse domains (Dubey et al., 2024). DeepSeek-R1-Distill-Llama-8B [^5] https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Llama-8B: An 8-billion-parameter distilled variant leveraging knowledge transfer from the 671B DeepSeek-R1 model, specialized in step-by-step mathematical reasoning and efficient deployment on consumer-grade hardware (Guo et al., 2025b).
Unless otherwise specified, all generative processes are conducted using the vLLM inference framework with a temperature setting of $T=0.5$. The system prompt used for response generation was: You are a helpful assistant. For the reasoning-oriented model, we first force the output of $\n\n</think>$ and then let the model continue generating, thereby implementing the “without CoT” setting.
@@
TruthfulQA offers a training set with 6.9k examples, which we use to fine-tune the GPT-3-6.7B model for evaluation purposes. Since OpenAI’s API no longer provides access to this model, we instead fine-tune the Llama-3.1-8B [^6] https://huggingface.co/meta-llama/Llama-3.1-8B model using LoRA (Hu et al., 2022) to compute the Truthfulness and Informativeness scores. We conduct fine-tuning on the LLaMA-Factory [^7] https://github.com/hiyouga/LLaMA-Factory platform with the following hyperparameters: learning rate = 5.0e-05, batch size = 32, epochs = 5, lora_rank = 8, and lora_alpha = 16.
The evaluation metrics are defined in Section 4.1. For Expected Calibration Error (ECE), we apply a normalization process to ensure predictions fall within the $[0,1]$ range. Specifically, we scale the data within the $3\%-97\%$ range to $[0,1]$, while outliers beyond this range are clipped to 0 or $1$. This normalization avoids distortions caused by extreme values. Verbalized Certainty is excluded from this analysis because its hallucination detection scores are confined to five distinct levels.